OpenStudy (baldymcgee6):

How to use the quadratic formula on (x^2+1)/(x-1)) =3^4

OpenStudy (baldymcgee6):

\[\frac{ x^2+1 }{ x-1 } =3^4\]

OpenStudy (baldymcgee6):

\[(x^2+1)/(x-1)) -3^4 = 0\]

OpenStudy (anonymous):

\[ x ^{2}+1 =3^{4}(x-1)\]

OpenStudy (baldymcgee6):

got it, forgot.. thanks again

OpenStudy (anonymous):

expand, group like terms, etc.